How much force will it take to lift a 10 pounds weight with a wheel and axle if the


axle is 2 inches in diameter and the wheel is 10 inches in diameter?

1
Expert's answer
2021-11-14T17:15:17-0500

"R_a\cdot W=R_w\cdot F\to F=R_a\cdot W/R_w=1\cdot10/5=2\ (lb)"